The diagonal of a square television screen measures 18 inches. What is the length of the sides?
Answer by Earlsdon(6294)

The diagonal of a square forms the hypotenuse of a right triangle.
If you know the length of the hypotenuse of a right triangle, you can find the length of the sides using the Pythagorean theorem.
c^2 = a^2 + b^2 In the square with sides, a and b, a = b and c is the diagonal.
c^2 = 2a^2 Divide both sides by 2.
a^2 = c^2/2 Take the square root of both sides.
a = c/sqrt(2)
a = 18/sqrt(2) = 18/1.4142... = 12.73 inches, approximately.